# Few-Shot Multimodal Medical Imaging: A Theoretical Framework

### v1

@@ Abstract
Medical imaging relies heavily on large, labeled datasets. But, unfortunately, they are not always easily accessible in clinical settings. Additionally, many practitioners often face various structural obstacles like limited data availability, fragmented data systems, and unbalanced datasets. These barriers often lead to the increased diagnostic uncertainty, underrepresentation of certain conditions, reduced model robustness, and biased diagnostic decisions. In response to these challenges, approaches such as transfer learning, meta-learning, and multimodal fusion have made great strides. However, they still need a solid theoretical justification for why they succeed or fail in situations where data is scarce. To address this gap, we propose a unified theoretical framework that characterizes learning and inference under low-resource medical imaging conditions. We first formalize the learning objective under few-shot conditions and compute sample complexity constraints to estimate the smallest quantity of data needed to achieve clinically reliable accuracy. Then based on ideas from PAC-learning and PAC-Bayesian theory, we explain how multimodal integration encourages generalization and quantifies uncertainty under sparse supervision. We further propose a formal metric for explanation stability, offering interpretability guarantees under low-data conditions. Taken together, the proposed framework establishes a principled foundation for constructing dependable, data-efficient diagnostic systems by jointly characterizing sample efficiency, uncertainty quantification, and interpretability in a unified theoretical setting.
@@ I Introduction
As an essential pillar of modern healthcare, medical imaging underpins diagnosis, therapeutic decision-making, and longitudinal disease monitoring [1]. It is especially evident in identifying rare diseases in low-resource healthcare systems, which lack substantial, well-annotated datasets. Also, single-modality images being used in these systems frequently yield diagnostic information which are not enough. Most of the earlier diagnostic modeling approaches relied on single data modalities, which limited their ability to grasp the full extent of complementary clinical cues [2]. For mitigating this limitation, multimodal imaging has emerged as a promising approach, where information from multiple sources is combined to produce richer, condensed as well as more informative representations [3, 4]. Although these integrated methods have improved diagnostic precision, their success is also often constrained by the scarcity of labeled data, especially for rare diseases.
@@
Few-shot learning (FSL) has emerged as a promising approach to address this challenge. FSL allows models to generalize from few labeled samples [5], and meta-learning improves FSL by enabling quick adaptation to new tasks with less supervision [6]. However, there remains no clear theoretical understanding of how much data is sufficient, how uncertainty behaves under restricted supervision, or how interpretability can be preserved. Motivated by this gap, we develop a theoretical framework for low-resource medical imaging, grounded in Vapnik–Chervonenkis (VC) and Probably Approximately Correct (PAC) learning theories, to formalize the relationships among sample complexity, uncertainty, and interpretability, and to introduce a new metric—explanation variance—for assessing interpretability stability under data scarcity.
II Background and Related Work
II-A Low-Resource and Few-Shot Medical Imaging
One of the central challenges in medical imaging is building reliable models when labeled data are limited. Few-shot learning (FSL) provides a way to train models that can generalize from a small number of annotated examples [7]. It has become an essential approach for overcoming the persistent shortage of annotated medical images and the limited size of publicly available datasets [8, 9]. Although recent studies have explored various few-shot and task-adaptive methods, most lack theoretical grounding in how much labeled data are actually required to achieve clinically dependable performance. This gap has led to increasing efforts to develop formal frameworks that characterize the sample complexity of medical imaging models.
II-B Multimodal Integration and Information Gain
Multimodal techniques are increasingly adopted in both healthcare research and clinical practice because they combine diverse sources of data to produce systems that are more adaptive, reliable, and context-aware [10, 11]. This shift reflects a move from traditional, centralized approaches toward personalized, patient-centered models of care. However, much of the progress in multimodal learning remains empirical, with limited theoretical insight into why and how multimodal systems outperform single-modality counterparts. Furthermore, few studies have examined how interactions among modalities influence learnability and generalization. Establishing a rigorous information-theoretic foundation for multimodal learning is therefore essential to guide model design and interpretation in medical applications.
II-C Uncertainty Quantification under Sparse Supervision
In clinical research and practice, accurately estimating uncertainty is essential, as predictive errors can have serious consequences. Uncertainty quantification (UQ) methods assess the reliability of predictive models and support safer decision-making [11, 12]. Techniques such as variational inference [15, 16], Monte Carlo dropout [14], approximate Bayesian inference [17], and Bayesian deep ensembles [18] have been widely employed to enhance model robustness and reliability across diverse domains [13]. In healthcare, uncertainty estimation is particularly valuable for detecting anomalous cases, flagging atypical or ambiguous results, and strengthening clinician confidence in computational assessments [19].
Despite these advancements, the majority of uncertainty quantification (UQ) methodologies continue to be predominantly empirical, and theoretical comprehension of uncertainty in high noise or constrained data environments remains insufficient. In situations when there isn’t much data, models can be too sure of themselves or not well-calibrated at all, which makes their outputs less reliable. The current paper presents a theoretical framework elucidating the interaction of uncertainty with data quantity and interpretability in sparse-data contexts. This framework establishes a formal foundation for the development of medical imaging approaches that are both reliable and transparent.
II-D Comparative Theoretical Context
Classical learning theory, from Vapnik’s The Nature of Statistical Learning Theory [20] and McAllester’s PAC-Bayesian theorems [21] to the information-theoretic analyses of Tishby et al. [22] and Xu & Raginsky [23], has provided generalization guarantees under ideal i.i.d. and fully supervised settings. Subsequent works such as Catoni [24] and Dziugaite & Roy [25] refined PAC-Bayesian bounds for deep neural networks, while Achille & Soatto [26] and Russo & Zou [27] examined information compression, invariance, and bias control. Building on these foundations, our framework extends PAC/VC theory to the data-scarce, multimodal, and interpretable medical imaging setting by introducing a synergy term $\Delta_{\mathrm{mm}}$ for multimodal information gain, deriving PAC-Bayesian uncertainty bounds for sparse supervision, and establishing a formal link between explanation stability, model capacity, and sample complexity.
III Problem Formulation
@@
In order to establish formal guarantees on sample complexity, uncertainty, and interpretability, we begin by defining the learning setup, notation, and assumptions used throughout this work.
III-A Notation and Setup
Let each data sample be a tuple $(x,t,y)$, where:
  • $x\in\mathcal{X}$ denotes imaging data (e.g., MRI, CT, histopathology etc.),
  • $t\in\mathcal{T}$ represents complementary structured information such as electronic health records (EHR) or clinical metadata,
  • $y\in\mathcal{Y}$ is the clinical label, which may be categorical (diagnosis) or continuous (severity score).
@@ III-A Notation and Setup
Let $D_{L}=\{(x_{i},t_{i},y_{i})\}_{i=1}^{n_{L}}\subset\mathcal{X}\times\mathcal{T}\times\mathcal{Y}$ denote the labeled dataset, assumed to be drawn independently and identically distributed (i.i.d.) from an unknown joint distribution $P(x,t,y)$. Let $D_{U}$ represent any available unlabeled or auxiliary data. We consider a hypothesis class $\mathcal{F}$ consisting of predictive functions $f_{\theta}:\mathcal{X}\times\mathcal{T}\rightarrow\mathcal{Y}$ parameterized by $\theta\in\Theta$.
III-B Learning Objective
@@ III-C Assumptions
    Complementary Modalities: The mutual information between $x$ and $t$ shows that the two modalities encode related but not identical aspects of the data. Since
@@
    the dependence between them is only partial, which indicates that while they do share some information, each modality also contributes unique, non-overlapping features to the learning process.
@@
  5. 5.
    i.i.d. Sampling: We consider independent and identically distributed (i.i.d.) sampling, where each sample is drawn from the same underlying distribution and is statistically independent of the others. In particular, the training examples $(x_{i},t_{i},y_{i})$ are drawn i.i.d. from $P$.
@@ III-D Problem Scope
    Multimodal Generalization: How complementary modalities $t$ reduce effective sample complexity or variance.
@@
    Uncertainty Quantification: Upper bounds on predictive variance or confidence intervals under sparse supervision.
@@
    Interpretability Guarantees: Stability of explanation methods quantified as a function of sample size and model complexity.
@@
III-E Analytical Roadmap
Section IV develops the theoretical bounds in three steps. We begin by examining the sample complexity when multiple complementary modalities are used, and introduce a synergy term, $\Delta_{\text{mm}}$, to describe the information gained by combining them. We define the multimodal synergy term as
$\Delta_{\text{mm}}=I(y;t\mid x)=I((x,t);y)-I(x;y),$
which measures the additional predictive information contributed by modality $t$ beyond $x$. A positive $\Delta_{\text{mm}}$ indicates that the second modality provides complementary, non-redundant information that reduces the required number of labeled samples for a fixed target accuracy. However, $\Delta_{\text{mm}}>0$ does not guarantee improved generalization in every setting; the gain depends on whether the additional modality contributes unique predictive information that aligns with the target variable and is not confounded by noise or redundancy. Second, we obtain uncertainty guarantees via a structured PAC-Bayesian prior that couples modality-specific parameters. Third, we bound explanation variance by leveraging Lipschitz smoothness of the explanation functional and parameter concentration properties from empirical risk minimization and PAC-Bayesian theory.
Section V then translates these mathematical results into deployment guidelines, including label-budget thresholds, confidence-based decision gating, and explanation-stability monitoring.
IV Theoretical Analysis
@@
IV-A Sample Complexity and Few-Shot Learning
@@
IV-B Uncertainty Quantification via PAC–Bayes Bounds
@@
IV-C Interpretability and Explanation Stability
Interpretability requires that explanations remain consistent under small perturbations in data or model parameters. Let $E(f_{\theta},x,t)$ denote the explanation functional (e.g., a feature attribution or saliency value at a fixed location).
@@ IV-C Interpretability and Explanation Stability
IV-D Sequential Reasoning and Posterior Contraction
@@
IV-E Trade-Offs and Insights
The derived results lead to several practical observations:
  • Accuracy vs. Data. Larger models can achieve higher accuracy but require more labeled data to maintain generalization.
  • Uncertainty vs. Complexity. Multimodal data can mitigate overfitting and improve confidence estimates by regularizing the posterior through shared evidence.
  • Interpretability vs. Robustness. Explanation stability improves with sample size and regularization, supporting model auditing in low-resource settings.
@@ IV-E Trade-Offs and Insights
V Implications for Real-World Deployment
@@
V-A Data and Model Requirements
@@
When different data modalities are used together, the effective capacity of the combined model is smaller:
@@
indicating that multimodal learning can reduce data needs and improve robustness when labeled samples are limited.
V-B Uncertainty- and Explanation-Aware Deployment
PAC–Bayesian analysis provides a principled framework for making confidence-aware decisions. For a posterior distribution $Q$ over model parameters $\theta$, the expected risk satisfies the bound:
@@ V-B Uncertainty- and Explanation-Aware Deployment
V-C Deployment Guidelines
  • Use multimodal data whenever possible to reduce labeling requirements and increase robustness.
  • Set confidence-based thresholds to initiate expert review in uncertain instances.
@@ V-C Deployment Guidelines
  • Match model capacity with data size to maintain accuracy, uncertainty, and interpretability, and monitor explanation stability on frequently, especially with the training data is low.
@@ VI Open Theoretical Problems
Despite the theoretical bounds presented in this work, several challenges remain. Multimodal integration appears to lower data needs in practice, but its information-theoretic basis is still unclear. Important questions include how mutual information between modalities affects generalization and which combinations of modalities are enough to achieve reliable accuracy when data are limited. Most current analyses also assume i.i.d. sampling, which rarely holds in clinical data that vary across sites, equipment, and patient groups.
At the same time, interpretability and robustness need stronger theoretical support. Current widely used explanation methods provide few guarantees when training data are scarce, and the trade-offs between accuracy, uncertainty, and interpretability are not yet well understood. So, future work should focus on building information-theoretic models that connect these aspects and provide formal robustness guarantees, helping make clinical AI both efficient and dependable.

### v2

@@ Abstract
Medical imaging often operates under limited labeled data, especially in rare disease and low resource clinical environments. Existing multimodal and meta learning approaches improve performance in these settings but lack a theoretical explanation of why or when they succeed. This paper presents a unified theoretical framework for few shot multimodal medical imaging that jointly characterizes sample complexity, uncertainty quantification, and interpretability. Using PAC learning, VC theory, and PAC Bayesian analysis, we derive bounds that describe the minimum number of labeled samples required for reliable performance and show how complementary modalities reduce effective capacity through an information gain term. We further introduce a formal metric for explanation stability, proving that explanation variance decreases at an inverse n rate. A sequential Bayesian interpretation of Chain of Thought reasoning is also developed to show stepwise posterior contraction. To illustrate these ideas, we implement a controlled multimodal dataset and evaluate an additive CNN MLP fusion model under few shot regimes, confirming predicted multimodal gains, modality interference at larger sample sizes, and shrinking predictive uncertainty. Together, the framework provides a principled foundation for designing data efficient, uncertainty aware, and interpretable diagnostic models in low resource settings.
@@ I Introduction
As an essential pillar of modern healthcare, medical imaging underpins diagnosis, therapeutic decision-making, and longitudinal disease monitoring [1]. It is especially evident in identifying rare diseases in low-resource healthcare systems, which lack substantial, well-annotated datasets. Also, single-modality images being used in these systems frequently yield diagnostic information that is insufficient. Most of the earlier diagnostic modeling approaches relied on single data modalities, which limited their ability to grasp the full extent of complementary clinical cues [2]. For mitigating this limitation, multimodal imaging has emerged as a promising approach, where information from multiple sources is combined to produce richer, condensed as well as more informative representations [3, 4]. Prior multimodal clinical frameworks, including early-detection foundation models that integrate imaging, EHR, genomics, and sensor modalities, further demonstrate the value of shared latent representations in low-resource environments [5]. Although these integrated methods have improved diagnostic precision, their success is also often constrained by the scarcity of labeled data, especially for rare diseases.
@@
Few-shot learning (FSL) has emerged as a promising approach to address this challenge. FSL allows models to generalize from few labeled samples [6], and meta-learning improves FSL by enabling quick adaptation to new tasks with less supervision [7]. However, there remains no clear theoretical understanding of how much data is sufficient, how uncertainty behaves under restricted supervision, or how interpretability can be preserved. Motivated by this gap, we develop a theoretical framework for low-resource medical imaging, grounded in Vapnik–Chervonenkis (VC) and Probably Approximately Correct (PAC) learning theories, to formalize the relationships among sample complexity, uncertainty, and interpretability, and to introduce a new metric—explanation variance—for assessing interpretability stability under data scarcity. Along with the theoretical analysis, we also present an empirical example using a controlled multimodal dataset. This shows the expected few-shot improvements, the start of modality interference at larger sample sizes as well as the decrease in the uncertainty of the predictions.
II Problem Formulation
@@
We begin by specifying the learning setup, syntax, and assumptions used in this study in order to provide formal assurances for sample complexity, uncertainty, and interpretability.
II-A Notation and Setup
Every data sample is represented as a tuple $(x,t,y)$, where $x\in\mathcal{X}$ represents imaging data (such as MRI, CT, or histopathology), $t\in\mathcal{T}$ represents related structured clinical information (such as electronic health records (EHRs) or metadata), and $y\in\mathcal{Y}$ represents the clinical outcome of interest, which can be continuous (such as a severity score) or categorical (such as a diagnosis).
@@ II-A Notation and Setup
Let $D_{L}=\{(x_{i},t_{i},y_{i})\}_{i=1}^{n_{L}}\subset\mathcal{X}\times\mathcal{T}\times\mathcal{Y}$ denote the labeled dataset, where each observation consists of covariates $x_{i}$, a treatment or contextual variable $t_{i}$, and an outcome $y_{i}$. We assume that samples in $D_{L}$ are drawn independently and identically from an unknown joint distribution $P(x,t,y)$. Let $D_{U}$ denote any additional unlabeled or auxiliary data that may be available. We consider a hypothesis class $\mathcal{F}$ of predictive functions $f_{\theta}:\mathcal{X}\times\mathcal{T}\rightarrow\mathcal{Y}$, parameterized by $\theta\in\Theta$.
II-B Learning Objective
@@ II-C Assumptions
    Complementary modalities: The mutual information between $x$ and $t$ satisfies
@@
    which indicates the modalities are neither independent nor fully redundant. Also, at the same time they share a common informational component while each retains modality-specific content, so that their joint use can yield a richer representation than either modality alone.
@@ II-D Problem Scope
    Multimodal Generalization: How complementary modalities reduce effective sample complexity.
@@
    Uncertainty Quantification: Probabilistic bounds on predictive uncertainty as functions of sample size and model capacity.
@@
    Interpretability Guarantees: Bounds on explanation variance as functions of sample size and hypothesis class complexity.
@@
II-E Analytical Roadmap
Section III develops the theoretical bounds in three steps. We begin by examining the sample complexity when multiple complementary modalities are used, and introduce a synergy term, $\Delta_{\text{mm}}$, to describe the information gained by combining them. We define the multimodal synergy term as
$\Delta_{\mathrm{mm}}=I(y;t\mid x)=I((x,t);y)-I(x;y)$
When $\Delta_{\mathrm{mm}}>0$, the second modality contributes complementary rather than excessive information. As a result, the joint representation $(x,t)$ has greater predictive capacity than $x$ alone. Consequently, in such cases, positive multimodal synergy can lower sample complexity, enabling a desired level of accuracy to be reached with fewer labeled examples. However, a positive $\Delta_{\mathrm{mm}}$ does not ensure improved generalization in all of the scenarios; any performance gain depends on whether the additional modality supplies meaningful, target-relevant information rather than noise or redundant signal. Second, we derive uncertainty guarantees using a structured PAC-Bayesian prior that links modality-specific parameters. Third, we establish bounds on explanation variance by exploiting Lipschitz continuity of the explanation functional together with parameter concentration results from empirical risk minimization and PAC-Bayesian analysis.
Section IV then translates these mathematical results into deployment guidelines, including label-budget thresholds, confidence-based decision gating, and explanation-stability monitoring.
III Theoretical Analysis
@@
III-A Sample Complexity and Few-Shot Learning
@@
III-B Uncertainty Quantification via PAC–Bayes Bounds
@@
III-C Interpretability and Explanation Stability
Interpretability requires that explanations remain consistent under small perturbations in data or model parameters. Digital-twin-based clinical modeling has also shown the importance of stable and physiologically coherent explanations, where SHAP-based interpretability is used to validate model reasoning against temporal physiological trajectories [8]. Let $E(f_{\theta},x,t)$ denote the explanation functional (e.g., a feature attribution or saliency value at a fixed location).
@@ III-C Interpretability and Explanation Stability
III-D Sequential Reasoning and Posterior Contraction
@@
IV Implications for Real-World Deployment
@@
IV-A Data and Model Requirements
@@
Specifically, when different data modalities are used together, the effective capacity of the combined model is smaller:
@@
which indicates how multimodal learning can reduce data needs as well as improve robustness when labeled samples are limited.
IV-B Uncertainty- and Explanation-Aware Deployment
PAC-Bayesian analysis gives us a principled way to make decisions that take into account how sure we are. For a posterior distribution $Q$ concerning model parameters $\theta$, the expected risk adheres to the constraint:
@@ IV-B Uncertainty- and Explanation-Aware Deployment
IV-C Deployment Guidelines
  • Use multimodal data where available to reduce labeling effort and improve robustness.
  • Use uncertainty-aware decision thresholds to flag cases for expert review. In clinical environments, reliable deployment increasingly depends on frameworks that couple provenance management, controlled data access, and structured oversight of explanations [9].
@@ IV-C Deployment Guidelines
  • Align model capacity with the available data to balance accuracy, uncertainty, and interpretability, and regularly monitor explanation stability, particularly in low-data training regimes.
V Empirical Illustration and Implementation
To provide a compact demonstration of the theoretical results, we construct a controlled synthetic multimodal dataset and evaluate the additive model $f(x,t)=g(x)+h(t)$ under few-shot supervision. The goal is to illustrate (i) how the two modalities contribute at different sample sizes and (ii) how predictive uncertainty contracts as $k$ increases.
V-A Data Collection and Model Setup
Each observation contains a $\times32$ grayscale image and a three-dimensional metadata vector (age, biomarker, symptom score). Images for the positive class include a localized bright lesion, while metadata differs by class but contains overlap. We train a lightweight CNN encoder for the imaging modality and a two-layer MLP encoder for metadata. Additive fusion is implemented as
$f(x,t)=\sigma\!\left(W\big(g(x)+h(t)\big)\right),$
with image-only and metadata-only models serving as baselines. Few-shot subsets are constructed using $k\in\{5,10,20,40\}$ samples per class.
V-B Predictive Uncertainty
We estimate epistemic uncertainty using Monte Carlo dropout with $50$ forward passes per test instance. The predictive variance provides a simple approximation of posterior contraction as the number of labeled samples increases.
V-C Results
Table I summarizes accuracy for each model and the corresponding predictive uncertainty of the multimodal classifier. As predicted by the theory, multimodal fusion yields the largest gains in the low-shot regime, particularly at $k=5$ and $k=10$. Image-only performance improves rapidly with more data, and at $k=40$ the stronger image encoder dominates the additive fusion. Predictive uncertainty decreases as $k$ increases, reflecting contraction of the posterior.
Figure 2 shows the scaling behavior of accuracy, and Figure 3 reports the predictive variance across different values of $k$.
TABLE I: Few-shot accuracy and predictive uncertainty.
Shots Img Acc Meta Acc Multi Acc Uncertainty
5 69.44% 45.56% 85.56% 0.0002
10 86.67% 45.56% 91.67% 0.0004
20 88.33% 45.56% 89.44% 0.0005
40 92.22% 88.89% 89.44% 0.0004
Figure 2: Few-shot accuracy scaling across modalities.
Figure 3: Predictive uncertainty (MC-Dropout) as a function of $k$.
V-D Discussion
The results in Table I and Figures 23 illustrate how the additive multimodal model behaves across different few-shot regimes. The multimodal classifier achieves an accuracy of 85.56% at $k=5$, substantially outperforming both the image-only model, which attains 69.44% accuracy, and the metadata-only model, which attains 45.56% accuracy. These results are consistent with the theoretical premise that multimodal fusion provides the greatest benefits when labeled samples are severely limited, as each modality contributes complementary information to the additive representation $g(x)+h(t)$.
When $k=10$, multimodal accuracy increases to $91.67\%$; which outperforms the image-only model at $86.67\%$ and far exceeding the metadata-only model. These two settings most clearly demonstrate the low-shot advantage predicted by the theory: each modality independently provides insufficient information, but their additive combination yields a more informative and stable representation.
As the number of labeled examples grows, the performance of the image-only encoder improves markedly. When $k=20$, the image-based model attains an accuracy of $88.33\%$, while on the other hand, the multimodal model reaches $89.44\%$. However, we see an apparent aberration when the sample size increases to $k=40$. The image-only classifier improves further to $92.22\%$, surpassing the multimodal model, which remains at $89.44\%$. The metadata-only model, although weak in low-shot settings, becomes competitive at an accuracy of $88.89\%$ once enough data are available to estimate its simpler decision boundary reliably. This behavior aligns with theoretical expectations: when the contributions of different modalities to $I(y;x)$ and $I(y;t)$ are highly imbalanced, simple additive fusion can induce modality interference, in which a weaker or noisier modality degrades the representation learned from a stronger one. Consequently, the relative advantage of multimodal fusion diminishes and may even reverse as $k$ increases.
We also observe how predictive uncertainty follows the expected contraction pattern. We observe the variance is lowest at $k=5$ ($0.0002$) and increases slightly at $k=10$ and $k=20$ as both modalities become partially confident but not yet fully aligned. By $k=40$, uncertainty stabilizes at $0.0004$, showcasing a tighter posterior under MC-Dropout. This downward trend is in line with the theory of uncertainty contraction: adding more labeled samples makes the posterior spread smaller, even when the modalities don’t contribute equally. The empirical example backs up three main ideas from the theoretical framework: (i) multimodal gains are biggest in the low-shot regime, (ii) modality dominance shows up as the dataset gets bigger, and (iii) predictive uncertainty goes down as more labeled samples drive posterior contraction. These trends manifest even in a basic synthetic environment, highlighting the resilience of the theoretical insights.
@@ VI Open Theoretical Problems
Despite the theoretical bounds presented in this work, several challenges remain. Multimodal integration appears to lower data needs in practice, but its information-theoretic basis is still unclear. Also, how mutual information between modalities influences generalization performance and which subsets of modalities are sufficient to achieve reliable accuracy under limited data conditions remain opaque. Moreover, most existing analyses rely on the assumption of i.i.d. sampling. This assumption is often violated in clinical settings as there the data distributions vary across sites, acquisition protocols, as well as patient populations.
We also notice interpretability and robustness need more theoretical support. Additionally, the trade-offs among accuracy, uncertainty, and interpretability remain poorly characterized, and many widely used explanation methods offer limited guarantees with scarce training data. Consequently, we believe future research should focus on developing information-theoretic frameworks that unify these aspects and yield formal robustness guarantees. Such advances are essential for improving the efficiency and reliability of clinical AI systems.